A boy rides a bike 8 miles. The trip would have taken 20 minutes less if he had ridden 2 miles per hour faster. find his actual speed.
Found 3 solutions by nerdybill, stanbon, mangopeeler07:
Answer by nerdybill(7384)

A boy rides a bike 8 miles. The trip would have taken 20 minutes less if he had ridden 2 miles per hour faster. find his actual speed.
.
20 minutes = 20/60 hours = 1/3 hour
.
Let x = actual speed
and y = actual time taken
.
xy = 8 (equation 1)
(x+2)(y-1/3) = 8 (equation 2)
.
solving equation 1 for y:
xy = 8
y = 8/x
.
substitute the above into equation 2 and solve for x:
(x+2)(y-1/3) = 8
(x+2)(8/x-1/3) = 8
multiply both sides by 3x:
(x+2)(24-x) = 24x
24x + 48 - x^2 - 2x = 24x
22x + 48 - x^2= 24x
0 = x^2 + 2x - 48
0 = (x+8)(x-6)
x = {-8, 6}
.
Since speed can't be negative, it must be the positive solution.
.
x = 6 mph
.

Answer by stanbon(75887)

A boy rides a bike 8 miles. The trip would have taken 20 minutes less if he had ridden 2 miles per hour faster. find his actual speed.
-----------------------------------
Original DATA:
distance = 8 miles ; rate = x mph ; time = d/r = 8/x hrs
------------------
Modified DATA:
distance = 8 miles ; rate = (x+2) mph ; time = d/r = 8/(x+2)
-------------------
EQUATION:
old time - new time = 20/60 = 1/3 hr
8/x + 8/(x+2) = 1/3
Multiply thru by 3x(x+2)
24(x+2) + 24x = x(x+2)
48x + 48 = x^2 + 2x
x^2 - 46x - 48 = 0
-------
x = [46 +- sqrt(46^2 -4*1*-48)]/2
Positive solution:
x = [46 + sqrt(2300)]/2
x = 47.02.. mph (actual speed)
================================
Cheers,
Stan H.

Answer by mangopeeler07(462)

time=distance/speed
t=8/s

t-1/3=8/s+2

Plug in t

8/s-1/3=8/s+2

Common denominator on the left side=3s
24/3s-s/3s=8/s+2

Subtract the numerators
24-s/3s=8/s+2

Cross multiply
24s=-s^2+26s+48

Subtract the 24s from both sides
0=-s^2+2s+48

Factor
0=(-s+8)(s+6)

Then you solve for s
s=-6 or 8

Speed can't be negative, so make them positive
s=6 or 8

Plug in 6.
t=8/6

t=4/3

4/3-1/3=8/6+2
It is true!

there fore the original speed was 6 mph.